OpenStudy (anonymous):

HELP!!!!! I don't understand the simplification of this problem can you explain it to me please. I think I'm just missing something or there is a missed step somewhere. I am putting in the comments the equation. How is the answer after the arrow gotten (I put the problem in the comments.

OpenStudy (anonymous):

\[\sqrt{2m}\left( \frac{ a }{ 2 } \right)\left( \sqrt{E} + \sqrt{E-V}\right) = n \pi \hbar\] \[\rightarrow E + E - V +2\sqrt{E \left( E-V \right)} = \frac{ 4 }{ 2m }\left( \frac{ n \pi \hbar }{ a } \right)^{2}\]